OpenStudy (anonymous):

solve the equation. express radicals in simplest form. (x+9)^2=12 i got -9+2sqrt3, -9-2sqrt3

OpenStudy (lalaly):

thats right

OpenStudy (anonymous):

x^2 + 18x + 81 = 12 x^2 + 18x + 69 = 0 x = -9+2sqrt3, -9-2sqrt3